OpenStudy (anonymous):

find the area enclosed by the x-axis and the curve x=5+e^t y= t-t^2

OpenStudy (anonymous):

range of t is?

OpenStudy (amistre64):

it might be easier to define t in terms of x; then sub that into y

OpenStudy (amistre64):

x=5+e^t x-5 = e^t ln|x-5| = t

OpenStudy (anonymous):

Well, you could convert that to some awful function of x, which would then give you \[\int_6^{5+e} f(x) dx\] but I'd much prefer to make the substitution into the variable t. Letting \[x = 5 + e^t, dx = e^t dt\] that gives \[\int_0^1 (t-t^2)e^t dt = 3-e\]

OpenStudy (amistre64):

y = t-t^2 roots on the xaxis at 0 = t(1-t); 0 and 1 ln|x-5| = 0, when x=5 ln|x-5| = 1, when x = e+5 ... such a nice number that is too Jems looks easier by far, but I got no idea how to go about conceiving of it :)

OpenStudy (amistre64):

6-5 = 1 ..... yeah

OpenStudy (amistre64):

int: ln|x-5| = (x-5) (ln|x-5| - 1) if i recall the parttern correctly

OpenStudy (anonymous):

In principle you can of course solve for t and plug it in to get a function of x. That would yield an integral \[\int \ln(x-5)-\ln^2(x-5) dx\] with bounds of six and 5+e corresponding to the points t=0 and t=1. All I'm saying is don't even bother making the substitution, because converting the above integrand back to a function of t just gives you \[\int (t-t^2) e^t dt\] with that extra exponential thrown in there because of the dx -> e^t dt substitution.

OpenStudy (anonymous):

Plugging the first integral into wolfram alpha yields the same thing, but I surely wouldn't want to do it by hand :)

OpenStudy (anonymous):

thank you everyone for your help :)